OpenStudy (anonymous):

What is the derivative of y=tan(ln(2x+1))

OpenStudy (tkhunny):

(d/dx) tan(x) = (sec(x))^2 (d/dx)ln(x) = 1/x Chain, chain, chain.....

OpenStudy (anonymous):

Thank you very much! I forgot that the derivative of tan(x) = (secx)^2

OpenStudy (tkhunny):

That's one you may have to look up a few times before it sticks. You can observe that tan(x) = sin(x)/cos(x) and practice your Quotient Rule if you want to demonstrate it. :-)